OpenStudy (anonymous):

I need to evaluate this integral..... x/(2x-1)^1/2 with an upper =5 and the lower =1. Can anyone Help?

OpenStudy (anonymous):

let's re-write it in a better form first: \[\int\limits_{1}^{5}x/\sqrt{2x-1}dx\] where you can take : \[ f' =({2x-1} )^{-1/2}\] g = x so now you must find g' and f in this case :) you can take the sqrt of (2x-1) up and solve using the following rule : \[\int\limits_{}^{}f'gdx = fg - \int\limits_{}^{}fg'dx\] give it a try ^_^

OpenStudy (anonymous):

substitute and solve :)

OpenStudy (anonymous):

How do you take the antiderivative of (2x-1)^-1/2???